1. A confined amount of gas at 100C has a pressure of 2.5 atm. What is its new pressure if the temperature is increased to 50oC?


1
Expert's answer
2021-06-25T05:41:04-0400

P1/T1 = P2/T2

2.5/373 = P2/323

0.009156 = P2/323

P2 = 0.009156×323

P2 = 2.96atm